OpenStudy (anonymous):

Can anyone help me with finding the equation of a line with the points (-3,-4) and (1,-2)?

OpenStudy (anonymous):

first find the slope

OpenStudy (anonymous):

y2-y1/x2-x1 ---> slope formula

OpenStudy (anonymous):

first you need to find the slope

OpenStudy (anonymous):

(-2-(-4))/1-(-3) =2/4 = 1/2 is the slope

OpenStudy (anonymous):

now use this formula to find the equation of a line y-y1 = m(x-x1) m represents the slope now just plug the point in to the equation

OpenStudy (anonymous):

which would be y-(-4) =(1/2)(x-(-3)) y+4 =(1/2)(x+3) y = 1/2x + 3/2 + 4 y = (1/2)x + 11/2